# Solving Inverse Problems with FLAIR

### v1

@@ 5.2 Experimental Results
In the case of image inpainting, our method produces high-quality reconstructions that fully leverage the power of the generative model and blend naturally into the surrounding context, avoiding degradations and artifacts that we observe in the baselines. In particular, FlowDPS tends to produce implausible textures in the inpainted regions, while FlowChef regularly fails to generate semantically consistent content at all. For single-image super-resolution, FLAIR consistently delivers the most perceptually convincing and realistic outputs. Notably, the FID scores remain low for both ×$\times 8$ and ×$\times 12$ magnification, indicating an effective usage of the generative prior to overcome the increasing ill-posedness. Again, FlowDPS suffers from blur and low texture quality, whereas FlowChef tends to lose semantic coherence. In motion deblurring, FLAIR also restores sharper and semantically more credible content than competing approaches, which often suffer from residual blur or inconsistent details. The boost in reconstruction quality is quantitatively reflected by all metrics, confirming that FLAIR reconstructs images with high fidelity. For further qualitative examples, please see subsection A.9.
@@
Table 1: Quantitative results with 50 NFE and $\sigma_{\nu}=0.01$.
@@
FFHQ 768$\times$768
ReSample 0.400 331 109 762 191 80.40033110976219180.400\,331\,109\,762\,191\,80.400 331 109 762 191 8 55.586 227 416 992 1955.5862274169921955.586\,227\,416\,992\,1955.586 227 416 992 19 0.814 689 397 811 889 60.81468939781188960.814\,689\,397\,811\,889\,60.814 689 397 811 889 6 26.367 465 972 900 3926.3674659729003926.367\,465\,972\,900\,3926.367 465 972 900 39 0.473 518 252 372 741 70.47351825237274170.473\,518\,252\,372\,741\,70.473 518 252 372 741 7 80.268 859 863 281 2580.2688598632812580.268\,859\,863\,281\,2580.268 859 863 281 25 0.785 562 992 095 947 30.78556299209594730.785\,562\,992\,095\,947\,30.785 562 992 095 947 3 25.469 425 201 416 01625.46942520141601625.469\,425\,201\,416\,01625.469 425 201 416 016 0.457 411 199 808 120 70.45741119980812070.457\,411\,199\,808\,120\,70.457 411 199 808 120 7 82.892 143 249 511 7282.8921432495117282.892\,143\,249\,511\,7282.892 143 249 511 72 0.787 788 689 136 505 10.78778868913650510.787\,788\,689\,136\,505\,10.787 788 689 136 505 1 25.446 083 068 847 65625.44608306884765625.446\,083\,068\,847\,65625.446 083 068 847 656 0.365 910 947 322 845 460.365910947322845460.365\,910\,947\,322\,845\,460.365 910 947 322 845 46 70.832 069 396 972 6670.8320693969726670.832\,069\,396\,972\,6670.832 069 396 972 66 0.827 21.827 827 453 613 2821.8278274536132821.827\,827\,453\,613\,2821.827 827 453 613 28
FlowDPS 0.374 372 541 904 449 460.374372541904449460.374\,372\,541\,904\,449\,460.374 372 541 904 449 46 40.299 484 252 929 6940.2994842529296940.299\,484\,252\,929\,6940.299 484 252 929 69 0.756 326 615 810 394 30.75632661581039430.756\,326\,615\,810\,394\,30.756 326 615 810 394 3 29.238 021 850 585 93829.23802185058593829.238\,021\,850\,585\,93829.238 021 850 585 938 0.413 043 200 969 696 040.413043200969696040.413\,043\,200\,969\,696\,040.413 043 200 969 696 04 47.244 354 248 046 87547.24435424804687547.244\,354\,248\,046\,87547.244 354 248 046 875 0.741 228 818 893 432 60.74122881889343260.741\,228\,818\,893\,432\,60.741 228 818 893 432 6 28.05 0.431 281 954 050 064 10.43128195405006410.431\,281\,954\,050\,064\,10.431 281 954 050 064 1 49.485 298 156 738 2849.4852981567382849.485\,298\,156\,738\,2849.485 298 156 738 28 0.731 617 748 737 335 20.73161774873733520.731\,617\,748\,737\,335\,20.731 617 748 737 335 2 27.637 281 417 846 6827.6372814178466827.637\,281\,417\,846\,6827.637 281 417 846 68 0.344 38.3 0.770 720 958 709 716 80.77072095870971680.770\,720\,958\,709\,716\,80.770 720 958 709 716 8 19.191 648 483 276 36719.19164848327636719.191\,648\,483\,276\,36719.191 648 483 276 367
RSD 0.391 473 561 525 344 850.391473561525344850.391\,473\,561\,525\,344\,850.391 473 561 525 344 85 51.667 427 062 988 2851.6674270629882851.667\,427\,062\,988\,2851.667 427 062 988 28 0.776 033 043 861 389 20.77603304386138920.776\,033\,043\,861\,389\,20.776 033 043 861 389 2 29.690 019 607 543 94529.69001960754394529.690\,019\,607\,543\,94529.690 019 607 543 945 0.462 387 174 367 904 660.462387174367904660.462\,387\,174\,367\,904\,660.462 387 174 367 904 66 71.734 550 476 074 2271.7345504760742271.734\,550\,476\,074\,2271.734 550 476 074 22 0.743 28.108 963 012 695 31228.10896301269531228.108\,963\,012\,695\,31228.108 963 012 695 312 0.457 823 425 531 387 330.457823425531387330.457\,823\,425\,531\,387\,330.457 823 425 531 387 33 77.319 137 573 242 1977.3191375732421977.319\,137\,573\,242\,1977.319 137 573 242 19 0.743 024 885 654 449 50.74302488565444950.743\,024\,885\,654\,449\,50.743 024 885 654 449 5 27.67 0.478 281 795 978 546 140.478281795978546140.478\,281\,795\,978\,546\,140.478 281 795 978 546 14 73.293 403 625 488 2873.2934036254882873.293\,403\,625\,488\,2873.293 403 625 488 28 0.735 569 655 895 233 20.73556965589523320.735\,569\,655\,895\,233\,20.735 569 655 895 233 2 21.97
FlowChef 0.341 30.5 0.760 072 886 943 817 10.76007288694381710.760\,072\,886\,943\,817\,10.760 072 886 943 817 1 28.424 997 329 711 91428.42499732971191428.424\,997\,329\,711\,91428.424 997 329 711 914 0.373 41.0 0.730 421 423 912 048 30.73042142391204830.730\,421\,423\,912\,048\,30.730 421 423 912 048 3 26.996 358 871 459 9626.9963588714599626.996\,358\,871\,459\,9626.996 358 871 459 96 0.406 35.0 0.716 119 468 212 127 70.71611946821212770.716\,119\,468\,212\,127\,70.716 119 468 212 127 7 25.809 129 714 965 8225.8091297149658225.809\,129\,714\,965\,8225.809 129 714 965 82 0.394 163 548 946 380 60.39416354894638060.394\,163\,548\,946\,380\,60.394 163 548 946 380 6 71.046 714 782 714 8471.0467147827148471.046\,714\,782\,714\,8471.046 714 782 714 84 0.780 103 147 029 876 70.78010314702987670.780\,103\,147\,029\,876\,70.780 103 147 029 876 7 18.184 221 267 700 19518.18422126770019518.184\,221\,267\,700\,19518.184 221 267 700 195
Ours 0.213 235 303 759 574 90.21323530375957490.213\,235\,303\,759\,574\,90.213 235 303 759 574 9 13.266 211 509 704 5913.2662115097045913.266\,211\,509\,704\,5913.266 211 509 704 59 0.777 29.54 0.270 564 854 145 050 050.270564854145050050.270\,564\,854\,145\,050\,050.270 564 854 145 050 05 16.201 229 095 458 98416.20122909545898416.201\,229\,095\,458\,98416.201 229 095 458 984 0.740 209 460 258 483 90.74020946025848390.740\,209\,460\,258\,483\,90.740 209 460 258 483 9 27.714 468 002 319 33627.71446800231933627.714\,468\,002\,319\,33627.714 468 002 319 336 0.236 155 018 210 411 070.236155018210411070.236\,155\,018\,210\,411\,070.236 155 018 210 411 07 10.732 836 723 327 63710.73283672332763710.732\,836\,723\,327\,63710.732 836 723 327 637 0.772 29.611 394 882 202 1529.6113948822021529.611\,394\,882\,202\,1529.611 394 882 202 15 0.183 834 627 270 698 550.183834627270698550.183\,834\,627\,270\,698\,550.183 834 627 270 698 55 8.741 712 570 190 438.741712570190438.741\,712\,570\,190\,438.741 712 570 190 43 0.827 794 373 035 430 90.82779437303543090.827\,794\,373\,035\,430\,90.827 794 373 035 430 9 23.693 054 199 218 7523.6930541992187523.693\,054\,199\,218\,7523.693 054 199 218 75
DIV2K 768$\times$768
ReSample $0.533$ 54.997 371 673 583 9854.9973716735839854.997\,371\,673\,583\,9854.997 371 673 583 98 $0.625$ $22.34$ 0.642 829 596 996 307 40.64282959699630740.642\,829\,596\,996\,307\,40.642 829 596 996 307 4 88.051 658 630 371 188.051658630371188.051\,658\,630\,371\,188.051 658 630 371 1 0.561 500 668 525 695 80.56150066852569580.561\,500\,668\,525\,695\,80.561 500 668 525 695 8 $20.85$ 0.555 853 188 037 872 30.55585318803787230.555\,853\,188\,037\,872\,30.555 853 188 037 872 3 79.654 029 846 191 479.654029846191479.654\,029\,846\,191\,479.654 029 846 191 4 0.617 21.792 427 062 988 2821.7924270629882821.792\,427\,062\,988\,2821.792 427 062 988 28 0.285 51.859 077 453 613 2851.8590774536132851.859\,077\,453\,613\,2851.859 077 453 613 28 0.796 22.676 401 138 305 66422.67640113830566422.676\,401\,138\,305\,66422.676 401 138 305 664
FlowDPS 0.476 54.160 160 064 697 26654.16016006469726654.160\,160\,064\,697\,26654.160 160 064 697 266 0.567 384 541 034 698 50.56738454103469850.567\,384\,541\,034\,698\,50.567 384 541 034 698 5 23.005 729 675 292 9723.0057296752929723.005\,729\,675\,292\,9723.005 729 675 292 97 0.547 238 528 728 485 10.54723852872848510.547\,238\,528\,728\,485\,10.547 238 528 728 485 1 67.424 385 070 800 7867.4243850708007867.424\,385\,070\,800\,7867.424 385 070 800 78 0.528 21.79 0.558 460 950 851 440 40.55846095085144040.558\,460\,950\,851\,440\,40.558 460 950 851 440 4 60.432 220 458 984 37560.43222045898437560.432\,220\,458\,984\,37560.432 220 458 984 375 0.535 993 039 608 001 70.53599303960800170.535\,993\,039\,608\,001\,70.535 993 039 608 001 7 21.877 470 016 479 49221.87747001647949221.877\,470\,016\,479\,49221.877 470 016 479 492 0.327 635 318 040 847 80.32763531804084780.327\,635\,318\,040\,847\,80.327 635 318 040 847 8 34.2 0.691 909 790 039 062 50.69190979003906250.691\,909\,790\,039\,062\,50.691 909 790 039 062 5 21.714 540 481 567 38321.71454048156738321.714\,540\,481\,567\,38321.714 540 481 567 383
RSD 0.538 948 655 128 4790.5389486551284790.538\,948\,655\,128\,4790.538 948 655 128 479 60.930 324 554 443 3660.9303245544433660.930\,324\,554\,443\,3660.930 324 554 443 36 0.591 312 229 633 331 30.59131222963333130.591\,312\,229\,633\,331\,30.591 312 229 633 331 3 23.445 989 608 764 6523.4459896087646523.445\,989\,608\,764\,6523.445 989 608 764 65 0.683 576 822 280 8830.6835768222808830.683\,576\,822\,280\,8830.683 576 822 280 883 95.726 249 694 824 2295.7262496948242295.726\,249\,694\,824\,2295.726 249 694 824 22 0.523 275 852 203 369 10.52327585220336910.523\,275\,852\,203\,369\,10.523 275 852 203 369 1 21.964 975 357 055 66421.96497535705566421.964\,975\,357\,055\,66421.964 975 357 055 664 0.637 668 132 781 982 40.63766813278198240.637\,668\,132\,781\,982\,40.637 668 132 781 982 4 97.646 888 732 910 1697.6468887329101697.646\,888\,732\,910\,1697.646 888 732 910 16 0.551 110 327 243 804 90.55111032724380490.551\,110\,327\,243\,804\,90.551 110 327 243 804 9 22.10 0.463 778 406 381 607 060.463778406381607060.463\,778\,406\,381\,607\,060.463 778 406 381 607 06 63.870 880 126 953 12563.87088012695312563.870\,880\,126\,953\,12563.870 880 126 953 125 0.677 523 434 162 139 90.67752343416213990.677\,523\,434\,162\,139\,90.677 523 434 162 139 9 23.23
FlowChef 0.490 322 709 083 557 130.490322709083557130.490\,322\,709\,083\,557\,130.490 322 709 083 557 13 48.6 0.539 492 428 302 764 90.53949242830276490.539\,492\,428\,302\,764\,90.539 492 428 302 764 9 21.842 645 645 141 621.842645645141621.842\,645\,645\,141\,621.842 645 645 141 6 0.525 55.3 0.492 154 121 398 925 80.49215412139892580.492\,154\,121\,398\,925\,80.492 154 121 398 925 8 20.521 999 359 130 8620.5219993591308620.521\,999\,359\,130\,8620.521 999 359 130 86 0.561 51.3 0.486 047 297 716 140 750.486047297716140750.486\,047\,297\,716\,140\,750.486 047 297 716 140 75 19.899 375 915 527 34419.89937591552734419.899\,375\,915\,527\,34419.899 375 915 527 344 0.488 857 418 298 721 30.48885741829872130.488\,857\,418\,298\,721\,30.488 857 418 298 721 3 86.286 514 282 226 5686.2865142822265686.286\,514\,282\,226\,5686.286 514 282 226 56 0.659 262 299 537 658 70.65926229953765870.659\,262\,299\,537\,658\,70.659 262 299 537 658 7 20.874 286 651 611 32820.87428665161132820.874\,286\,651\,611\,32820.874 286 651 611 328
Ours 0.352 998 495 101 928 70.35299849510192870.352\,998\,495\,101\,928\,70.352 998 495 101 928 7 26.456 972 122 192 38326.45697212219238326.456\,972\,122\,192\,38326.456 972 122 192 383 0.607 23.30 0.421 490 550 041 198 730.421490550041198730.421\,490\,550\,041\,198\,730.421 490 550 041 198 73 32.118 663 787 841 832.118663787841832.118\,663\,787\,841\,832.118 663 787 841 8 0.525 365 948 677 0630.5253659486770630.525\,365\,948\,677\,0630.525 365 948 677 063 21.385 475 158 691 40621.38547515869140621.385\,475\,158\,691\,40621.385 475 158 691 406 0.314 553 380 012 512 20.31455338001251220.314\,553\,380\,012\,512\,20.314 553 380 012 512 2 21.111 848 831 176 75821.11184883117675821.111\,848\,831\,176\,75821.111 848 831 176 758 0.652 780 890 464 782 70.65278089046478270.652\,780\,890\,464\,782\,70.652 780 890 464 782 7 24.444 496 154 785 15624.44449615478515624.444\,496\,154\,785\,15624.444 496 154 785 156 0.162 881 195 545 196 530.162881195545196530.162\,881\,195\,545\,196\,530.162 881 195 545 196 53 11.026 673 316 955 56611.02667331695556611.026\,673\,316\,955\,56611.026 673 316 955 566 0.814 652 204 513 549 80.81465220451354980.814\,652\,204\,513\,549\,80.814 652 204 513 549 8 23.752 101 898 193 3623.7521018981933623.752\,101\,898\,193\,3623.752 101 898 193 36
@@
Posterior Variance. To demonstrate that FLAIR does not suffer from mode collapse, we assess the posterior variance $\operatorname*{Var}[x|z]$ for the task of ×$\times 12$ Super Resolution, by drawing 32 samples for a fixed observation $z$ and computing their pixel-wise variance. We conduct that experiment for our FLAIR approach, for RSD with repulsive term, and for FlowDPS kim2025flowdps . The example in Figure 3 illustrates that FLAIR has the highest sample diversity, which is also reflected in the corresponding variance maps. Notably, the sample variance is concentrated in regions with high-frequency textures. This indicates that our method reliably reconstructs the posterior, whose low-frequency part is, in the super-resolution setting, tightly constrained by the likelihood term.
@@
Figure 3: Zoomed-in reconstructions for x12 Super Resolution. We show posterior samples (col. 1–4) of FLAIR, FlowDPS, and RSD, posterior mean and standard deviation (over 32 samples, col. 5,6). 00 Refer to caption $0.16$
@@
Editing. Beyond image restoration, we observe that our method also performs remarkably well for text-based image editing, by simply presenting suitable target prompts during inpainting. Figure 4 illustrates a variety of edited images generated from the same photograph with the help of the depicted masks and prompts.

### v2

@@ 5.2 Experimental Results
In the case of image inpainting, our method produces high-quality reconstructions that fully leverage the power of the generative model and blend naturally into the surrounding context, avoiding degradations and artifacts that we observe in the baselines. In particular, FlowDPS tends to produce implausible textures in the inpainted regions, while FlowChef regularly fails to generate semantically consistent content at all.
For single-image super-resolution, FLAIR consistently delivers the most perceptually convincing and realistic outputs. Notably, the FID scores remain low for both $\times 8$ and $\times 12$ magnification, indicating an effective usage of the generative prior to overcome the increasing ill-posedness. Again, FlowDPS suffers from blur and low texture quality, whereas FlowChef tends to lose semantic coherence.
In motion deblurring, FLAIR also restores sharper and semantically more credible content than competing approaches, which often suffer from residual blur or inconsistent details. The boost in reconstruction quality is quantitatively reflected by all metrics, confirming that FLAIR reconstructs images with high fidelity. For further qualitative examples, see subsection A.13.
Refer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurringRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truthBox inpaintingRefer to captionobservation yyRefer to captionFlowDPSRefer to captionFlowChefRefer to captionFLAIRRefer to captionground truth $\times 12$ super-resolutionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionMotion deblurring
@@
Table 1: Quantitative results with 50 NFE and $\sigma_{\nu}=0.5\%$.
@@
FFHQ $\times768$
ReSample 0.400 331 109 762 191 80.400\,331\,109\,762\,191\,8 55.586 227 416 992 1955.586\,227\,416\,992\,19 0.814 689 397 811 889 60.814\,689\,397\,811\,889\,6 26.367 465 972 900 3926.367\,465\,972\,900\,39 0.473 518 252 372 741 70.473\,518\,252\,372\,741\,7 80.268 859 863 281 2580.268\,859\,863\,281\,25 0.785 562 992 095 947 30.785\,562\,992\,095\,947\,3 25.469 425 201 416 01625.469\,425\,201\,416\,016 0.457 411 199 808 120 70.457\,411\,199\,808\,120\,7 82.892 143 249 511 7282.892\,143\,249\,511\,72 0.787 788 689 136 505 10.787\,788\,689\,136\,505\,1 25.446 083 068 847 65625.446\,083\,068\,847\,656 0.365 910 947 322 845 460.365\,910\,947\,322\,845\,46 70.832 069 396 972 6670.832\,069\,396\,972\,66 0.827 21.827 827 453 613 2821.827\,827\,453\,613\,28
FlowDPS 0.374 372 541 904 449 460.374\,372\,541\,904\,449\,46 $38.528$ 0.756 326 615 810 394 30.756\,326\,615\,810\,394\,3 29.238 021 850 585 93829.238\,021\,850\,585\,938 0.413 043 200 969 696 040.413\,043\,200\,969\,696\,04 44.0 0.741 228 818 893 432 60.741\,228\,818\,893\,432\,6 28.05 0.431 281 954 050 064 10.431\,281\,954\,050\,064\,1 $54.300$ 0.731 617 748 737 335 20.731\,617\,748\,737\,335\,2 27.637 281 417 846 6827.637\,281\,417\,846\,68 0.344 42.5 0.770 720 958 709 716 80.770\,720\,958\,709\,716\,8 19.191 648 483 276 36719.191\,648\,483\,276\,367
RSD 0.391 473 561 525 344 850.391\,473\,561\,525\,344\,85 51.667 427 062 988 2851.667\,427\,062\,988\,28 0.776 033 043 861 389 20.776\,033\,043\,861\,389\,2 29.690 019 607 543 94529.690\,019\,607\,543\,945 0.462 387 174 367 904 660.462\,387\,174\,367\,904\,66 71.734 550 476 074 2271.734\,550\,476\,074\,22 0.743 28.108 963 012 695 31228.108\,963\,012\,695\,312 0.457 823 425 531 387 330.457\,823\,425\,531\,387\,33 77.319 137 573 242 1977.319\,137\,573\,242\,19 0.743 024 885 654 449 50.743\,024\,885\,654\,449\,5 27.67 0.478 281 795 978 546 140.478\,281\,795\,978\,546\,14 73.293 403 625 488 2873.293\,403\,625\,488\,28 0.735 569 655 895 233 20.735\,569\,655\,895\,233\,2 21.97
FlowChef 0.341 30.5 0.760 072 886 943 817 10.760\,072\,886\,943\,817\,1 28.424 997 329 711 91428.424\,997\,329\,711\,914 0.373 $46.5$ 0.730 421 423 912 048 30.730\,421\,423\,912\,048\,3 26.996 358 871 459 9626.996\,358\,871\,459\,96 0.406 40.2 0.716 119 468 212 127 70.716\,119\,468\,212\,127\,7 25.809 129 714 965 8225.809\,129\,714\,965\,82 0.394 163 548 946 380 60.394\,163\,548\,946\,380\,6 $69.837$ 0.780 103 147 029 876 70.780\,103\,147\,029\,876\,7 18.184 221 267 700 19518.184\,221\,267\,700\,195
Ours 0.213 235 303 759 574 90.213\,235\,303\,759\,574\,9 13.266 211 509 704 5913.266\,211\,509\,704\,59 0.777 29.54 0.270 564 854 145 050 050.270\,564\,854\,145\,050\,05 16.201 229 095 458 98416.201\,229\,095\,458\,984 0.740 209 460 258 483 90.740\,209\,460\,258\,483\,9 27.714 468 002 319 33627.714\,468\,002\,319\,336 0.236 155 018 210 411 070.236\,155\,018\,210\,411\,07 10.732 836 723 327 63710.732\,836\,723\,327\,637 0.772 29.611 394 882 202 1529.611\,394\,882\,202\,15 0.183 834 627 270 698 550.183\,834\,627\,270\,698\,55 8.741 712 570 190 438.741\,712\,570\,190\,43 0.827 794 373 035 430 90.827\,794\,373\,035\,430\,9 23.693 054 199 218 7523.693\,054\,199\,218\,75
DIV2K $\times768$
ReSample $0.533$ 54.997 371 673 583 9854.997\,371\,673\,583\,98 $0.625$ $22.34$ 0.642 829 596 996 307 40.642\,829\,596\,996\,307\,4 88.051 658 630 371 188.051\,658\,630\,371\,1 0.561 500 668 525 695 80.561\,500\,668\,525\,695\,8 $20.85$ 0.555 853 188 037 872 30.555\,853\,188\,037\,872\,3 79.654 029 846 191 479.654\,029\,846\,191\,4 0.617 21.792 427 062 988 2821.792\,427\,062\,988\,28 0.285 51.859 077 453 613 2851.859\,077\,453\,613\,28 0.796 22.676 401 138 305 66422.676\,401\,138\,305\,664
FlowDPS 0.476 $44.414$ 0.567 384 541 034 698 50.567\,384\,541\,034\,698\,5 23.005 729 675 292 9723.005\,729\,675\,292\,97 0.547 238 528 728 485 10.547\,238\,528\,728\,485\,1 $54.046$ 0.528 21.79 0.558 460 950 851 440 40.558\,460\,950\,851\,440\,4 $65.460$ 0.535 993 039 608 001 70.535\,993\,039\,608\,001\,7 21.877 470 016 479 49221.877\,470\,016\,479\,492 0.327 635 318 040 847 80.327\,635\,318\,040\,847\,8 29.2 0.691 909 790 039 062 50.691\,909\,790\,039\,062\,5 21.714 540 481 567 38321.714\,540\,481\,567\,383
RSD 0.538 948 655 128 4790.538\,948\,655\,128\,479 60.930 324 554 443 3660.930\,324\,554\,443\,36 0.591 312 229 633 331 30.591\,312\,229\,633\,331\,3 23.445 989 608 764 6523.445\,989\,608\,764\,65 0.683 576 822 280 8830.683\,576\,822\,280\,883 95.726 249 694 824 2295.726\,249\,694\,824\,22 0.523 275 852 203 369 10.523\,275\,852\,203\,369\,1 21.964 975 357 055 66421.964\,975\,357\,055\,664 0.637 668 132 781 982 40.637\,668\,132\,781\,982\,4 97.646 888 732 910 1697.646\,888\,732\,910\,16 0.551 110 327 243 804 90.551\,110\,327\,243\,804\,9 22.10 0.463 778 406 381 607 060.463\,778\,406\,381\,607\,06 63.870 880 126 953 12563.870\,880\,126\,953\,125 0.677 523 434 162 139 90.677\,523\,434\,162\,139\,9 23.23
FlowChef 0.490 322 709 083 557 130.490\,322\,709\,083\,557\,13 36.5 0.539 492 428 302 764 90.539\,492\,428\,302\,764\,9 21.842 645 645 141 621.842\,645\,645\,141\,6 0.525 43.8 0.492 154 121 398 925 80.492\,154\,121\,398\,925\,8 20.521 999 359 130 8620.521\,999\,359\,130\,86 0.561 49.6 0.486 047 297 716 140 750.486\,047\,297\,716\,140\,75 19.899 375 915 527 34419.899\,375\,915\,527\,344 0.488 857 418 298 721 30.488\,857\,418\,298\,721\,3 $58.339$ 0.659 262 299 537 658 70.659\,262\,299\,537\,658\,7 20.874 286 651 611 32820.874\,286\,651\,611\,328
Ours 0.352 998 495 101 928 70.352\,998\,495\,101\,928\,7 26.456 972 122 192 38326.456\,972\,122\,192\,383 0.607 23.30 0.421 490 550 041 198 730.421\,490\,550\,041\,198\,73 32.118 663 787 841 832.118\,663\,787\,841\,8 0.525 365 948 677 0630.525\,365\,948\,677\,063 21.385 475 158 691 40621.385\,475\,158\,691\,406 0.314 553 380 012 512 20.314\,553\,380\,012\,512\,2 21.111 848 831 176 75821.111\,848\,831\,176\,758 0.652 780 890 464 782 70.652\,780\,890\,464\,782\,7 24.444 496 154 785 15624.444\,496\,154\,785\,156 0.162 881 195 545 196 530.162\,881\,195\,545\,196\,53 11.026 673 316 955 56611.026\,673\,316\,955\,566 0.814 652 204 513 549 80.814\,652\,204\,513\,549\,8 23.752 101 898 193 3623.752\,101\,898\,193\,36
@@
Posterior Variance. To demonstrate that FLAIR does not suffer from mode collapse, we assess the posterior variance $\operatorname*{Var}[x|y]$ for the task of $\times 12$ Super Resolution, by drawing 32 samples for a fixed observation $y$ and computing their pixel-wise variance. We conduct that experiment for our FLAIR approach, for RSD with repulsive term, and for FlowDPS kim2025flowdps . The example in Figure 3 illustrates that FLAIR has the highest sample diversity, which is also reflected in the corresponding variance maps. Notably, the sample variance is concentrated in regions with high-frequency textures. This indicates that our method reliably reconstructs the posterior, whose low-frequency part is, in the super-resolution setting, tightly constrained by the likelihood term.
@@
Figure 3: Zoomed-in reconstructions for x12 Super Resolution. We show posterior samples (col. 1–4) of FLAIR, FlowDPS, and RSD, posterior mean and standard deviation (over 32 samples, col. 5,6). 0 Refer to caption $0.16$
@@
Editing. Beyond image restoration, we observe that our method also performs remarkably well for text-based image editing, simply presenting suitable target prompts during inpainting. Figure 4 illustrates a variety of edited images generated from the same photograph with the help of the depicted masks and prompts.
Refer to captionOriginal image Refer to caption"man wearing aviator glasses."Refer to caption"man wearing a wizard hat"Refer to caption"man with a Mike Tyson facial tattoo"Refer to caption"clown with a red nose,red makeup and a ruff"